OpenStudy (anonymous):

I really don't understand these questions. Read, and keep rereading and just not getting it. Please help !!! Collect like terms 2x – 5x 2x2 + 8x2 x – 5x 16x3 – 12x3

OpenStudy (abb0t):

(2-5)x = -3x Do you see what I did there? I basically subtracted the two numbers and then just multiplied it by x. That's what you're doing here.

OpenStudy (abb0t):

Same thing with the 3\(^{rd}\) one. (1-5)x = -4x now, you do the other two.

OpenStudy (anonymous):

Oh ok.. Now I understand.. Thank you very much, made it alot easier !

OpenStudy (skullpatrol):

This uses the distributive property: $$\Huge ba - ca = (b-c)a$$ $$\Huge 2x - 5x = (2-5)x$$ for a=x , b=2 , c=5